the complement of an angle is 13 more than 6 times the angle. find the value of the component
Answer by macston(5194)

x=the angle
.
90-x=6x+13
77=7x
11=x
90-x=90-11=79
ANSWER: The angle is 11 degrees and the complement is 79 degrees